a wheel on a game show is being pulled band starts to rotate with initial angular speed of 1.5 rad/s . it comes to rest after rotating through 3/4 of a turn . calculate the torque exerted on the wheel if the mass and the radius of the wheel is 6.5 kg and 0.7m respectively .
1
Expert's answer
2020-10-05T10:53:32-0400

A torque τ acting through an angle ∆θ does a work W is given by the equation

W = -τ∆θ

The kinemativ equation of angular velocity and angular acceleration is given by the equation

"\u03c9^2=\u03c9_0^2 + 2\u03b1\u2206\u03b8"

The wheel angular acceleration can be found using the kinematics equation as

"α = \frac{ω^2-ω_0^2}{2∆θ}"

"\u03c9_0" is initial angular speed and ω is final angular speed and ∆θ is angular displacement.

Since wheel come to rest after rotating through ¾ or 0.75 of a turn, its final angular velocity becomes zero (ω = 0).

In each turn wheel rotates rad hence the total change in displacement after come to rest can be expressed as

"∆θ = (0.75)(\frac{2π \;rad}{rev})"

Consider the wheel has a form of disk then moment of inertia of disk is given by the equation

"I = \frac{1}{2}mr^2"

The relation between torque and moment of inertia can be expressed as

"\u03c4 = I\u03b1"

Here negative sign indicates torque exerted on the wheel to stop

"τ = (\frac{1}{2}mr^2)(\frac{ω^2-ω_0^2}{2∆θ})"

"τ = -\frac{1}{4}(\frac{mr^2ω_0^2}{∆θ})"

Substitute 6.5 kg for mass of the disk (m), 0.7 m for radius of the disk (r), 1.5 rad/s for "\u03c9_0" and (0.75)"(\frac{2π \;rad}{rev})" for ∆θ in the above equation and solve for torque.

"τ = -\frac{1}{4}(\frac{(6.5\;kg)(0.7\;m)^2(1.5 \;rad/s)^2}{0.75(2π \;rad/rev)})"

"τ = -\frac{1}{4} (1.52\;N m)"

"τ = 0.38\;N m"